OpenStudy (anonymous):

The volume of a cylinder is 289 cm3 and the height is 21 cm. Find the diameter. Round the answer to the nearest whole number. A. 2 cm B. 3 cm C. 6 cm D. 4 cm

OpenStudy (anonymous):

im lost

OpenStudy (markrijckenberg):

Volume cylinder = 289 = pi * r(adius) * radius * height = pi * r * r * 21 Then 289 = 3.1415 * r * r * 21 Then 289 / 3.1415 = r * r * 21 Then 91.99 = r * r * 21 Then 91.99 / 21 = r * r Then 4.38 = r * r Then r = square root of 4.38 Then r ≃ 2.09 Then d = diameter = radius * 2 = 2.09 * 2 ≃ 4

OpenStudy (anonymous):

thanx